OpenStudy (anonymous):

What is the equation, in standard form, of a horizontal ellipse centered at (11, -6) with a major axis of length 16 and a minor axis of length 12? (1 points) Please don't just give me the answer

OpenStudy (anonymous):

\[\frac{(x-11)^2}{8^2}+\frac{(y+6)^2}{6^2}=1 \] standard form of ellipse.. since centre at 11,-6 x-11 and y+6.. 2a=dist. of major axis (X axis in thsi case)=8 nad 2b=6

OpenStudy (anonymous):

Would you mind checking my work for me because i solved the question and i got the wrong answer i want to know what my mistake was..

OpenStudy (anonymous):

What are the foci of the ellipse given by the equation (x-2)^2/ 36 + y-8^2/144 i did h =+/- a k 2 +/- c 8 c^2 = 64 - 4 c='s sqrt 60 but my options are 2 (+/-) 6 sqrt 3, 8 2 (+/-) 6 sqrt 5 , 8 2, 8(+/-) 6 sqrt 3 2,8 (+/-) 6 sqrt 5

OpenStudy (anonymous):

um i got 8+-6sqrt3